# Representability of the direct sum of uniform $q$-matroids

### v1

@@ 2. The direct sum of uniform q-matroids
Combining this with the assumption that $r+s+t=\dim_{\mathbb{F}_{q}}(I)>\rho_{\mathcal{S}}(I)$, we obtain
$\rho_{\mathcal{S}_{[t]\setminus[g]}}(X_{2})<s+t+r-k_{[a]}\leq k-k_{[a]}=k_{[t]\setminus[a]}.$
In addition, $\rho_{\mathcal{S}_{[t]\setminus[a]}}(X_{2})<s+t+r-k_{[a]}\leq s+t=\dim_{\mathbb{F}_{q}}(X_{2})$. These two last observations contradict (8), since they imply $\rho_{\mathcal{S}_{[t]\setminus[a]}}(X_{2})<\min\{k_{[t]\setminus[a]},\dim_{\mathbb{F}_{q}}(X_{2})\}$. ∎
@@ 3. A representation of the direct sum of uniform q-matroids
Then, $\mathcal{S}:=\mathcal{S}_{1}\oplus\ldots\oplus\mathcal{S}_{t}$ is a $(\Lambda_{k-1,\mathbf{k}},k-1)$-evasive $[n,k]_{q^{m}/q}$ system, with $n=n_{1}+\ldots+n_{t}$, $k=k_{1}+\ldots+k_{t}$ and $\mathbf{k}=(k_{1},\ldots,k_{t})$.
@@
and the vector $v^{q}:=(v_{1}^{q},\ldots,v_{\ell}^{q})$. Since $k_{i}<n_{i}$, then $\mathcal{S}_{i}$ is an $[n_{i},k_{i}]_{q^{m}/q}$ system, for each $i\in[t]$. Thus, $\mathcal{S}$ is an $[n,k]_{q^{m}/q}$ system. Let us prove that $\operatorname{wt}_{\mathcal{S}}(H)\leq k-1$ for each hyperplane $H$ not containing any of the spaces $\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$, for $i\in[t]$. Observe that such hyperplane will be of the form $H=(a_{1,1},\ldots,a_{1,k_{1}},a_{2,1},\ldots,a_{2,k_{2}},\ldots,a_{t,1},\ldots,a_{t,k_{t}})^{\perp}$, for some nonzero vectors $a_{i}=(a_{i,1},\ldots,a_{i,k_{i}})\in\mathbb{F}_{q^{m}}^{k_{i}}$, for $i\in[t]$. Equivalently, we will prove that
@@
We will show this using double induction. We first start our induction on $t$.
If $t=1$, then $|T|\leq q^{k_{1}-1}$, since $\mathcal{S}_{1}$ is a $(k_{1}-1)$-scattered $[n_{1},k_{1}]_{q^{m}/q}$ system.
@@
$T\subseteq T^{\prime}:=\left\{\left(-\sum_{r=2}^{t}a_{r}(x_{r}),x_{2},\ldots,x_{t}\right)\in\mathbb{F}_{q^{n_{2}}}\times\ldots\times\mathbb{F}_{q^{n_{t}}}\colon\sum_{r=2}^{t}d_{r}(x_{r})=0\right\}.$
By induction hypothesis, we have $|T^{\prime}|\leq q^{k_{2}+\ldots+(k_{t}+1)-1}=q^{k_{1}+\ldots+k_{t}-1}$.
@@
Note that (12) is an affine equation, so its number of solution in $\mathbb{F}_{q^{n_{1}}}\times\ldots\times\mathbb{F}_{q^{n_{s}}}$ is either 00 or equal to the cardinality of
@@
Let us assume that (9) holds for every $k_{1}^{\prime}<k_{1}$ and every $k_{2}^{\prime},\ldots,k_{t}^{\prime}\in\mathbb{N}$. Observe that, if $a_{1,k_{1}}=0$, then $a_{1}(x_{1})=a_{1}^{\prime}(x_{1})$, where $b_{1}=(a_{1,1},\ldots,a_{1,k_{1}-1})$, and
@@
$\begin{cases}-x_{1}^{q^{k_{1}-1}}-b_{1}(x_{1})=\displaystyle\sum\limits_{r=2}^{t}a_{r}(x_{r}),\\ -x_{1}^{q^{k_{1}-1}}-b_{1}^{q^{h}}(x_{1})=\displaystyle\sum\limits_{r=2}^{t-1}a_{r}^{q^{h}}(x_{r})+a_{t}^{q^{h}}(x_{r}^{q}),\end{cases}$
@@
$\begin{cases}\displaystyle\sum\limits_{r=s+1}^{t}d_{r}(x_{r})=0,\\ \displaystyle\sum\limits_{r=1}^{t}a_{r}(x_{r})=0,\end{cases}$ (14)
@@
$\sum_{r=1}^{s}a_{r}(x_{s})=-\sum_{r=s+1}^{t}a_{r}(\bar{x}_{r}).$ (15)
Note that (15) is an affine equation, so its number of solution in $\mathbb{F}_{q^{n_{1}}}\times\ldots\times\mathbb{F}_{q^{n_{s}}}$ is either 00 or equal to the cardinality of
$T^{\prime}:=\left\{(x_{1},\ldots,x_{s})\in\mathbb{F}_{q^{n_{1}}}\times\ldots\times\mathbb{F}_{q^{n_{s}}}\colon\sum_{r=1}^{s}a_{r}(x_{r})=0\right\}.$
@@
$q^{k_{s+1}+\ldots+(k_{t}+1)-1}q^{k_{1}+k_{2}+\ldots+k_{s}-1}=q^{k_{1}+k_{2}+\ldots+k_{t}-1}.$
@@
is representable. In particular, it is represantable over every field $\mathbb{F}_{q^{m}}$ such that $m=m_{1}\cdot\ldots\cdot m_{t}$ for any $m_{1},\ldots,m_{t}$ satisfying $\gcd(m_{i},m_{j})=1$ for each $i\neq j$ and $m_{i}\geq n_{i}$ for $i\in[t]$.
@@
such that $\mathcal{T}:=\mathcal{T}_{1}\oplus\ldots\oplus\mathcal{T}_{t}$ is an $[m_{1}+\ldots+m_{t},k]_{q^{m}/q}$ system which is $(\Lambda_{k-1,\mathbf{k}},k-1)$-evasive. Then, by choosing any $\mathbb{F}_{q}$-subspace $V_{i}\subseteq\mathbb{F}_{q^{m_{i}}}$, with $\dim_{\mathbb{F}_{q}}(V_{i})=n_{i}$ for $i\in[t]$, and defining
@@
we obtain that $\mathcal{S}:=\mathcal{S}_{1}\oplus\ldots\oplus\mathcal{S}_{t}$ is an $[n,k]_{q^{m}/q}$ system which is a $(\Lambda_{k-1,\mathbf{k}},k-1)$-evasive, with $n=n_{1}+\ldots+n_{t}$. Using Theorem 2.4 we conclude. ∎
@@ 4. The particular case of uniform q-matroids of rank 1
We have seen in Theorem 3.2 that the direct sum of uniform $q$-matroids is always representable. However, such a result does not characterize over which extension field $\mathbb{F}_{q^{m}}$ we can have a representation. In general, this is a difficult question. In this section, we will focus on the direct sum of $t$ many uniform $q$-matroids of rank $1$. We will first give some necessary conditions on $m$ and then we will give some sufficient conditions on $m$ ensuring that the direct sum of two uniform $q$-matroids of rank $1$ is representable.
@@ 4.1. Necessary conditions for the representability
For $i\in[t]$, let $n_{i}\geq 2$, and let $\mathcal{S}_{i}$ be an $[n_{i},1]_{q^{m}/q}$ system. Assume that $\mathcal{S}_{1}\oplus\ldots\oplus\mathcal{S}_{t}$ is $(\Lambda_{1,(1,\ldots,1)},1)$-evasive. Then, $n_{i}\leq\frac{m}{2}$ for each $i\in[t]$.
@@ 4.2. Construction for two summands over small extension fields
Theorem 2.4 shows that establishing the $\mathbb{F}_{q^{m}}$-representability of the direct sum $\mathcal{U}_{1,n_{1}}(q)\oplus\mathcal{U}_{1,n_{2}}(q)$ is equivalent to find an $[n_{1},1]_{q^{m}/q}$ system $\mathcal{S}_{1}$ and an $[n_{2},1]_{q^{m}/q}$ system $\mathcal{S}_{2}$ such that $\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is a $(\Lambda_{1,(1,1)},1)$-evasive $[n_{1}+n_{2},2]_{q^{m}/q}$ system. So, in the following we will exhibit examples of such subspaces. To this aim, we recall the following characterization of $(\Lambda_{1,(1,1)},1)$-evasiveness.
@@
Let $A,B\subseteq\mathbb{F}_{q^{m}}$ be two $\mathbb{F}_{q}$-subsapces with $\dim A=n_{1}$, $\dim B=n_{2}$. Then, $A\oplus B$ is $(\Lambda_{1,(1,1)},1)$-evasive if and only if for each choice of $a_{1},a_{2}\in A\setminus\{0\}$, $b_{1},b_{2}\in B\setminus\{0\}$ satisfying $a_{1}b_{2}=a_{2}b_{1}$, it must hold that $a_{1}=\lambda a_{2}$, $b_{1}=\lambda b_{2}$ for some $\lambda\in\mathbb{F}_{q}$. Equivalently, $A\oplus B$ is $(\Lambda_{1,(1,1)},1)$-evasive if and only if for every $a_{1},a_{2}\in A\setminus\{0\}$, $b_{1},b_{2}\in B\setminus\{0\}$ for which there exists $\lambda\in\mathbb{F}_{q^{m}}$ such that $(a_{1},a_{2})=\lambda(b_{1},b_{2})$, we have that $\lambda\in\mathbb{F}_{q}$.
@@
$\mathcal{S}_{1}$ =$=\{f(\gamma)\colon f\in\mathbb{F}_{q}[x]_{<n_{1}}\},$
$\mathcal{S}_{2}$ =$=\{g(\gamma^{n_{1}})\colon g\in\mathbb{F}_{q}[x]_{<n_{2}}\}.$
Then, $\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is a $(\Lambda_{1,(1,1)},1)$-evasive $[n_{1}+n_{2},2]_{q^{m}/q}$ system.
@@
In particular, since $\deg h\leq\max\{n_{1}\deg g_{1}+\deg f_{2},n_{1}\deg g_{2}+\deg f_{1}\}\leq n_{1}(n_{2}-1)+n_{1}-1<n_{1}n_{2}\leq m$, $h(\gamma)=0$ implies that all the coefficients of the polynomial $h(x)$ are 00. Thus, for each $i_{1}\in\{0,\ldots,n_{1}\}$, $i_{2}\in\{0,\ldots,n_{2}\}$ we must have $f_{1,i_{2}}g_{2,i_{1}}-f_{2,i_{2}}g_{1,i_{1}}=0$. Together with the assumption of $g_{1},g_{2}$ being monic, it readily follows that this implies $f_{1}=f_{2}$ and $g_{1}=g_{2}$, from which we conclude. ∎
@@
Then, $\mathcal{S}:=U\oplus\mathcal{S}_{2}$ with $\mathcal{S}_{2}=\{b+\xi b^{q}\colon b\in V\}$ is a $(\Lambda_{1,(1,1)},1)$-evasive $[n_{1}+n_{2},2]_{q^{m}/q}$ system.
@@
Let $r$ be a proper divisor of $m$ and let $\xi\in\mathbb{F}_{q^{m}}\setminus\mathbb{F}_{q^{r}}$. Let $\mathcal{S}_{1}:=\mathbb{F}_{q^{r}}$ and $\mathcal{S}_{2}:=\{a+\xi a^{q}\colon a\in\mathbb{F}_{q^{r}}\}$. Then, $\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is a $(\Lambda_{1,(1,1)},1)$-evasive $[2r,2]_{q^{m}/q}$ system.
@@
Let $q=p^{h}$ and $m=p^{r}$, and for each $i\in\{0,\ldots,p^{r}\}$, let $\mathcal{F}_{i}:=\ker p_{i}$, where $p_{i}$ is defined as in (16). Then:
@@
    $\mathcal{F}_{i}\cdot\mathcal{F}_{j}\subseteq\mathcal{F}_{i+j-1}$ for each $i,j$ such that $i+j\leq n+1$.
@@
On the other hand, an element $\beta\in V\setminus\{0\}$ belongs to $\xi V\setminus\{0\}$ if and only if $\xi=\beta\gamma^{-1}$ for some $\gamma\in V\setminus\{0\}$. Thus, each element in $V\setminus\{0\}$ is counted exactly $|V|-1=q^{\ell}-1$ times, and hence
@@
from which we can derive $12\ell\geq m+$ 1, contradicting the hypothesis. ∎

### v2

@@ 2. The direct sum of uniform q-matroids
Combining this with the assumption that $r+s+d=\dim_{\mathbb{F}_{q}}(I)>\rho_{\mathcal{S}}(I)$, we obtain
$\rho_{\mathcal{S}_{[t]\setminus[a]}}(X_{2})<s+d+r-k_{[a]}\leq k-k_{[a]}=k_{[t]\setminus[a]}.$
In addition, $\rho_{\mathcal{S}_{[t]\setminus[a]}}(X_{2})<s+d+r-k_{[a]}\leq s+d=\dim_{\mathbb{F}_{q}}(X_{2})$. These two last observations contradict (9), since they imply $\rho_{\mathcal{S}_{[t]\setminus[a]}}(X_{2})<\min\{k_{[t]\setminus[a]},\dim_{\mathbb{F}_{q}}(X_{2})\}$. ∎
Example 2.5.
Let $\mathcal{M}=\mathcal{U}_{1,2}(q)$. Then $\mathcal{M}$ is representable over $\mathbb{F}_{q^{2}}$, while it was shown in [19, Theorem 3.4] that $\mathcal{M}\oplus\mathcal{M}$ is representable over $\mathbb{F}_{q^{m}}$ if and only if $m\geq 4$. Let $q=2$, $m\geq 4$, and let $\mathbb{F}_{2^{4}}=\mathbb{F}_{2}(\alpha)$. We now show using Theorem 2.4 that a representation of $\mathcal{M}\oplus\mathcal{M}$ is given by the $4,2]_{2^{4}/2}$ system
$\mathcal{S}=\{(a,b)\in\mathbb{F}_{2^{m}}^{2}\;:\;a\in\langle 1,\alpha\rangle_{\mathbb{F}_{2}},b\in\langle 1,\alpha^{2}\rangle_{\mathbb{F}_{2}}\}\leq\mathbb{F}_{2^{m}}^{4}.$
Let $(a_{1},b_{1}),(a_{2},b_{2})\in\mathcal{S}$, with $a_{1},a_{2},b_{1},b_{2}\neq 0$, and assume that $(a_{1},b_{1})=\gamma(a_{2},b_{2})$ for some $\gamma\in\mathbb{F}_{2^{m}}$. Then $a_{1}b_{2}=a_{2}b_{1}$. By definition of $\mathcal{S}$, for each $i\in\{1,2\}$ we can write
$a_{i}=\lambda_{i,1}+\lambda_{i,2}\alpha,\quad b_{i}=\mu_{i,1}+\mu_{i,2}\alpha^{2},$
for some $\lambda_{i,j},\mu_{i,j}\in\mathbb{F}_{2}$, which implies
$(\lambda_{1,1}+\lambda_{1,2}\alpha)(\mu_{2,1}+\mu_{2,2}\alpha^{2})=(\lambda_{2,1}+\lambda_{2,2}\alpha)(\mu_{1,1}+\mu_{1,2}\alpha^{2}).$ (10)
Since $m\geq 4$ we have that $1,\alpha,\alpha^{2},\alpha^{3}$ are $\mathbb{F}_{2}$-linearly independent, and by equating their coefficients in (10), we obtain
$\begin{cases}\lambda_{1,1}\mu_{2,1}=\lambda_{2,1}\mu_{1,1},\\ \lambda_{1,2}\mu_{2,1}=\lambda_{2,2}\mu_{1,1},\\ \lambda_{1,1}\mu_{2,2}=\lambda_{2,1}\mu_{1,2},\\ \lambda_{1,2}\mu_{2,2}=\lambda_{2,2}\mu_{1,2}.\end{cases}$
Using the further assumption that $(\lambda_{i,1},\lambda_{i,2})\neq(0,0)$ and $(\mu_{i,1},\mu_{i,2})\neq(0,0)$ for $i\in\{1,2\}$ the only possible solutions satisfy
$\begin{cases}\lambda_{1,j}=\lambda_{2,j}&\mbox{ for }j\in\{1,2\},\\ \mu_{1,j}=\mu_{2,j}&\mbox{ for }j\in\{1,2\}.\end{cases}$
In other words, we get $\gamma=1$ and $(a_{1},b_{1})=(a_{2},b_{2})$. This shows that, for every $(a,b)\in\mathcal{S}$ with $a,b\neq 0$, we have
$\dim_{\mathbb{F}_{2}}(\mathcal{S}\cap\langle(a,b)\rangle_{\mathbb{F}_{2^{4}}})=1,$
and hence $\mathcal{S}$ is $(\Lambda_{(1,1),1},1)$-evasive. Thus, $\mathcal{S}$ is an $\mathbb{F}_{2^{4}}$-representation of $\mathcal{M}\oplus\mathcal{M}$. We will see a much more general version of this constructive proof for the representability of $\mathcal{U}_{1,n_{1}}(q)\oplus\mathcal{U}_{1,n_{2}}(q)$ in Proposition 4.6.
@@ 3. A representation of the direct sum of uniform q-matroids
Then, $\mathcal{S}:=\mathcal{S}_{1}\oplus\ldots\oplus\mathcal{S}_{t}$ is a $(\Lambda_{k-1,\mathbf{k}},k-1)$-evasive $n,k]_{q^{m}/q}$ system, with $n=n_{1}+\ldots+n_{t}$, $k=k_{1}+\ldots+k_{t}$ and $\mathbf{k}=(k_{1},\ldots,k_{t})$.
@@
and the vector $v^{q}:=(v_{1}^{q},\ldots,v_{\ell}^{q})$. Since $k_{i}<n_{i}$, then $\mathcal{S}_{i}$ is an $n_{i},k_{i}]_{q^{m}/q}$ system, for each $i\in[t]$. Thus, $\mathcal{S}$ is an $n,k]_{q^{m}/q}$ system. Let us prove that $\operatorname{wt}_{\mathcal{S}}(H)\leq k-1$ for each hyperplane $H$ not containing any of the spaces $\iota_{i}(\mathbb{F}_{q^{m}}^{{k_{i}}})$, for $i\in[t]$. Observe that such hyperplane will be of the form $H=(a_{1,1},\ldots,a_{1,k_{1}},a_{2,1},\ldots,a_{2,k_{2}},\ldots,a_{t,1},\ldots,a_{t,k_{t}})^{\perp}$, for some nonzero vectors $a_{i}=(a_{i,1},\ldots,a_{i,k_{i}})\in\mathbb{F}_{q^{m}}^{k_{i}}$, for $i\in[t]$. Equivalently, we will prove that
@@
We will show this using double induction on $t$ and $k_{1}$. We first start our induction on $t$.
If $t=1$, then $|T|\leq q^{k_{1}-1}$, since $\mathcal{S}_{1}$ is a $(k_{1}-1)$-scattered $n_{1},k_{1}]_{q^{m}/q}$ system.
@@
$T\subseteq T^{\prime}:=\left\{\left(-\sum_{r=2}^{t}a_{r}(x_{r}),x_{2},\ldots,x_{t}\right)\in\mathbb{F}_{q^{n_{1}}}\times\ldots\times\mathbb{F}_{q^{n_{t}}}\colon\sum_{r=2}^{t}d_{r}(x_{r})=0\right\}.$
Since $T^{\prime}$ has the same cardinality of the set
$\left\{(x_{2},\ldots,x_{t})\in\mathbb{F}_{q^{n_{2}}}\times\ldots\times\mathbb{F}_{q^{n_{t}}}\colon\sum_{r=2}^{t}d_{r}(x_{r})=0\right\},$
which only has t1t-1 components, we can use induction hypothesis on $t^{\prime}=t-1$, and derive that $|T^{\prime}|\leq q^{k_{2}+\ldots+(k_{t}+1)-1}=q^{k_{1}+\ldots+k_{t}-1}$.
@@
Note that (14) is an affine equation, so its number of solution in $\mathbb{F}_{q^{n_{1}}}\times\ldots\times\mathbb{F}_{q^{n_{s}}}$ is either 0 or equal to the cardinality of
@@
This concludes the base case $k_{1}=1$. Let us assume that (11) holds for every $k_{1}^{\prime}<k_{1}$ and every $k_{2}^{\prime},\ldots,k_{t}^{\prime}\in\mathbb{N}$. Observe that, if $a_{1,k_{1}}=0$, then $a_{1}(x_{1})=b_{1}(x_{1})$, where $b_{1}=(a_{1,1},\ldots,a_{1,k_{1}-1})$, and
@@
$\begin{cases}-x_{1}^{q^{k_{1}-1}}-b_{1}(x_{1})=\displaystyle\sum\limits_{r=2}^{t}a_{r}(x_{r}),\\ -x_{1}^{q^{k_{1}-1}}-b_{1}^{q^{h}}(x_{1})=\displaystyle\sum\limits_{r=2}^{t-1}a_{r}^{q^{h}}(x_{r})+a_{t}^{q^{h}}(x_{t}^{q}),\end{cases}$
@@
$\begin{cases}\displaystyle\sum\limits_{r=s+1}^{t}d_{r}(x_{r})=0,\\ \displaystyle\sum\limits_{r=1}^{t}d_{r}(x_{r})=0,\end{cases}$ (16)
@@
$\sum_{r=1}^{s}d_{r}(x_{r})=-\sum_{r=s+1}^{t}d_{r}(\bar{x}_{r}).$ (17)
Note that (17) is an affine equation, so its number of solution in $\mathbb{F}_{q^{n_{1}}}\times\ldots\times\mathbb{F}_{q^{n_{s}}}$ is either 0 or equal to the cardinality of
$T^{\prime}:=\left\{(x_{1},\ldots,x_{s})\in\mathbb{F}_{q^{n_{1}}}\times\ldots\times\mathbb{F}_{q^{n_{s}}}\colon\sum_{r=1}^{s}d_{r}(x_{r})=0\right\}.$
@@
$q^{k_{s+1}+\ldots+(k_{t}+1)-1}q^{k_{1}+k_{2}+\ldots+k_{s}-1}=q^{k_{1}+k_{2}+\ldots+k_{t}-1},$
concluding the proof. ∎
Example 3.2.
Let $t=3$ and $n_{1}=5$, $n_{2}=3$ and $n_{3}=2$, and consider the direct sum
$\mathcal{M}=\mathcal{U}_{2,5}(q)\oplus\mathcal{U}_{2,3}(q)\oplus\mathcal{U}_{1,2}(q).$
for any prime power $q$. Theorem 3.1 states that $\mathcal{M}$ is representable over $\mathbb{F}_{q^{m}}$ with $m=n_{1}n_{2}n_{3}=30$, and a representation is given by $(\mathcal{S},\rho_{\mathcal{S}})$ where $\mathcal{S}$ is the $10,5]_{q^{30}/q}$ system
$\mathcal{S}=\left\{(a,a^{q},b,b^{q},c)\,:\,a\in\mathbb{F}_{q^{5}},b\in\mathbb{F}_{q^{3}},c\in\mathbb{F}_{q^{2}}\right\}.$
@@
is representable. In particular, it is representable over every field $\mathbb{F}_{q^{m}}$ such that $m=m_{1}\cdot\ldots\cdot m_{t}$ for any $m_{1},\ldots,m_{t}$ satisfying $\gcd(m_{i},m_{j})=1$ for each $i\neq j$ and $m_{i}\geq n_{i}$ for $i\in[t]$.
@@
such that $\mathcal{T}:=\mathcal{T}_{1}\oplus\ldots\oplus\mathcal{T}_{t}$ is an $m_{1}+\ldots+m_{t},k]_{q^{m}/q}$ system which is $(\Lambda_{k-1,\mathbf{k}},k-1)$-evasive. Then, by choosing any $\mathbb{F}_{q}$-subspace $V_{i}\subseteq\mathbb{F}_{q^{m_{i}}}$, with $\dim_{\mathbb{F}_{q}}(V_{i})=n_{i}$ for $i\in[t]$, and defining
@@
we obtain that $\mathcal{S}:=\mathcal{S}_{1}\oplus\ldots\oplus\mathcal{S}_{t}$ is an $n,k]_{q^{m}/q}$ system which is a $(\Lambda_{k-1,\mathbf{k}},k-1)$-evasive, with $n=n_{1}+\ldots+n_{t}$. Using Theorem 2.4 we conclude. ∎
Example 3.4.
Let $t=3$ and $n_{1}=n_{2}=n_{3}=4$, and consider the direct sum
$\mathcal{M}=\mathcal{U}_{2,4}(q)\oplus\mathcal{U}_{2,4}(q)\oplus\mathcal{U}_{2,4}(q).$
for any prime power $q$. We consider three extension degrees $m_{1},m_{2},m_{3}$ such that $m_{i}\geq n_{i}$ for $i\in[3]$ and $\gcd(m_{i},m_{j})=1$ for each $1\leq i<j\leq 3$. A possible choice is $m_{1}=4$, $m_{2}=5$, $m_{3}=7$. Theorem 3.3 states that $\mathcal{M}$ is representable over $\mathbb{F}_{q^{m}}$ with $m=m_{1}m_{2}m_{3}=140$. In particular, following its proof, we can also construct the $12,6]_{q^{140}/q}$ system $\mathcal{S}$ such that $\mathcal{M}\cong(\mathcal{S},\rho_{\mathcal{S}})$. It is enough to fix three $4$-dimensional $\mathbb{F}_{q}$-subspaces $V_{1}=\mathbb{F}_{q^{4}}$, $V_{2}\subseteq\mathbb{F}_{q^{5}}$, $V_{3}\subseteq\mathbb{F}_{q^{7}}$, and take
$\mathcal{S}:=\left\{(a_{1},a_{1}^{q},a_{2},a_{2}^{q},a_{3},a_{3}^{q})\,:\,a_{i}\in V_{i},i\in[3]\right\}.$
@@ 4. The particular case of uniform q-matroids of rank 1
We have seen in Theorem 3.3 that the direct sum of uniform $q$-matroids is always representable. However, such a result does not characterize over which extension field $\mathbb{F}_{q^{m}}$ we can have a representation. In general, this is a difficult question. In this section, we will focus on the direct sum of $t$ many uniform $q$-matroids of rank $1$. We will first give some necessary conditions on $m$ and then we will give some sufficient conditions on $m$ ensuring that the direct sum of two uniform $q$-matroids of rank $1$ is representable over $\mathbb{F}_{q^{m}}$.
@@ 4.1. Necessary conditions for the representability
For $i\in[t]$, let $n_{i}\geq 2$, and let $\mathcal{S}_{i}$ be an $n_{i},1]_{q^{m}/q}$ system. Assume that $\mathcal{S}_{1}\oplus\ldots\oplus\mathcal{S}_{t}$ is $(\Lambda_{1,(1,\ldots,1)},1)$-evasive. Then, $n_{i}\leq\frac{m}{2}$ for each $i\in[t]$.
@@ 4.2. Construction for two summands over small extension fields
Theorem 2.4 shows that establishing the $\mathbb{F}_{q^{m}}$-representability of the direct sum $\mathcal{U}_{1,n_{1}}(q)\oplus\mathcal{U}_{1,n_{2}}(q)$ is equivalent to finding an $n_{1},1]_{q^{m}/q}$ system $\mathcal{S}_{1}$ and an $n_{2},1]_{q^{m}/q}$ system $\mathcal{S}_{2}$ such that $\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is a $(\Lambda_{1,(1,1)},1)$-evasive $n_{1}+n_{2},2]_{q^{m}/q}$ system. So, in the following we will exhibit examples of such subspaces. To this aim, we recall the following characterization of $(\Lambda_{1,(1,1)},1)$-evasiveness.
@@
Let $A,B\subseteq\mathbb{F}_{q^{m}}$ be two $\mathbb{F}_{q}$-subspaces with $\dim A=n_{1}$, $\dim B=n_{2}$. Then, $A\oplus B$ is $(\Lambda_{1,(1,1)},1)$-evasive if and only if for each choice of $a_{1},a_{2}\in A\setminus\{0\}$, $b_{1},b_{2}\in B\setminus\{0\}$ satisfying $a_{1}b_{2}=a_{2}b_{1}$, it must hold that $a_{1}=\lambda a_{2}$, $b_{1}=\lambda b_{2}$ for some $\lambda\in\mathbb{F}_{q}$. Equivalently, $A\oplus B$ is $(\Lambda_{1,(1,1)},1)$-evasive if and only if for every $a_{1},a_{2}\in A\setminus\{0\}$, $b_{1},b_{2}\in B\setminus\{0\}$ for which there exists $\lambda\in\mathbb{F}_{q^{m}}$ such that $(a_{1},a_{2})=\lambda(b_{1},b_{2})$, we have that $\lambda\in\mathbb{F}_{q}$.
@@
$\mathcal{S}_{1}$ $=\{f(\gamma)\colon f\in\mathbb{F}_{q}[x]_{<n_{1}}\},$
$\mathcal{S}_{2}$ $=\{g(\gamma^{n_{1}})\colon g\in\mathbb{F}_{q}[x]_{<n_{2}}\}.$
Then, $\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is a $(\Lambda_{1,(1,1)},1)$-evasive $n_{1}+n_{2},2]_{q^{m}/q}$ system.
@@
In particular, since $\deg h\leq\max\{n_{1}\deg g_{1}+\deg f_{2},n_{1}\deg g_{2}+\deg f_{1}\}\leq n_{1}(n_{2}-1)+n_{1}-1<n_{1}n_{2}\leq m$, $h(\gamma)=0$ implies that all the coefficients of the polynomial $h(x)$ are 0. Thus, for each $i_{1}\in\{0,\ldots,n_{1}\}$, $i_{2}\in\{0,\ldots,n_{2}\}$ we must have $f_{1,i_{2}}g_{2,i_{1}}-f_{2,i_{2}}g_{1,i_{1}}=0$. Together with the assumption of $g_{1},g_{2}$ being monic, it readily follows that this implies $f_{1}=f_{2}$ and $g_{1}=g_{2}$, from which we conclude. ∎
@@
Then, $\mathcal{S}:=U\oplus\mathcal{S}_{2}$ with $\mathcal{S}_{2}=\{b+\xi b^{q}\colon b\in V\}$ is a $(\Lambda_{1,(1,1)},1)$-evasive $n_{1}+n_{2},2]_{q^{m}/q}$ system.
@@
Let $r$ be a proper divisor of $m$ and let $\xi\in\mathbb{F}_{q^{m}}\setminus\mathbb{F}_{q^{r}}$. Let $\mathcal{S}_{1}:=\mathbb{F}_{q^{r}}$ and $\mathcal{S}_{2}:=\{a+\xi a^{q}\colon a\in\mathbb{F}_{q^{r}}\}$. Then, $\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is a $(\Lambda_{1,(1,1)},1)$-evasive $2r,2]_{q^{m}/q}$ system.
@@
Let $q=p^{h}$ and $m=p^{r}$, and for each $i\in\{0,\ldots,p^{r}\}$, let $\mathcal{F}_{i}:=\ker p_{i}\subseteq\mathbb{F}_{q^{m}}$, where $p_{i}$ is defined as in (18). Then:
@@
    $\mathcal{F}_{i}\cdot\mathcal{F}_{j}\subseteq\mathcal{F}_{i+j-1}$ for each $i,j$ such that $i+j\leq m+1$.
@@
On the other hand, an element $\beta\in V\setminus\{0\}$ belongs to $\xi V\setminus\{0\}$ if and only if $\xi=\beta\gamma^{-1}$ for some $\gamma\in V\setminus\{0\}$. Thus, each element in $V\setminus\{0\}$ is counted exactly $|V|-1=q^{\ell}-1$ times in the sum $\sum_{\xi\in\mathbb{F}_{q^{m}}^{*}}|(V\cap\xi V)\setminus\{0\}|$, and hence
@@
from which we can derive $2\ell\geq m+1$, contradicting the hypothesis. ∎